# HiFo-Prompt: Prompting with Hindsight and Foresight for LLM-based Automatic Heuristic Design

### v1

@@ A.3 Knowledge-Augmented Evolutionary Computation
To extend the capabilities of evolutionary algorithms beyond simple adaptive control, a prominent research direction involves incorporating an explicit knowledge component. Cultural Algorithms [28, 45, 65] provide a classic framework for this idea, decoupling the evolutionary system into two interacting spaces: a population space containing candidate solutions $P^{(t)}$ and a belief space $K_{t}$ serving as a repository of experiential knowledge [27]. These two spaces co-evolve through a dual-inheritance communication protocol.
@@
Concurrently, the successes of the population are fed back into the belief space. An acceptance function $\mathcal{A}$ identifies and extracts potentially valuable experiences, $\mathcal{A}(P^{(t)})$, from the current population. These are then used by a knowledge update function $\mathcal{U}_{K}$ to update the belief space:
@@ B.1 Traveling Salesman Problem
The Traveling Salesman Problem (TSP) [47, 66] is a canonical NP-hard combinatorial optimization problem. Given a set of $N$ cities and the distances between each pair, the objective is to find the shortest possible tour that visits each city exactly once before returning to the starting city [67].
@@
While fast, the myopic nature of such simple rules often leads to suboptimal solutions.
Our approach, HIFO-Prompt, introduces a new paradigm for solving combinatorial optimization problems by shifting from traditional, fixed construction heuristics to dynamically generated, instance-specific policies. We retain the foundational step-by-step framework of construction methods in which a solution is built incrementally. However, we fundamentally revolutionize the core decision-making logic. Instead of relying on a universal, handcrafted rule (e.g., Nearest Neighbor), which is often myopic and susceptible to poor initial choices leading to local optima, our method leverages a LLM to synthesize a sophisticated decision-making function on-the-fly for each problem instance. At every construction step $t$, the LLM acts as a reasoning engine. It is provided with the complete state of the problem, including the current partial tour $S_{t}$, the set of unvisited candidate nodes $U_{t}$, the global distance matrix $C$, and the tour’s origin node to facilitate reasoning about the final closing cost [9, 25].
The advantage of HIFO-Prompt lies in its ability to offload the complex, knowledge-intensive task of heuristic design to the LLM. Through carefully engineered prompts, we guide the model not merely to select a node but to generate a computational policy that embodies advanced strategic principles. The synthesized policy can perform non-trivial reasoning, such as conducting evaluations of multi-step consequences to look beyond immediate greedy gains. It can be guided to implement mechanisms analogous to maintaining a belief over a set of promising candidates, exploring their short-term implications before committing to a single path. Furthermore, the generated function can incorporate stochastic simulations or rollouts to approximate the long-term value of different choices, mimicking the exploration capabilities of advanced search algorithms.
@@
Improvement heuristics, such as local search, start with a complete tour and iteratively refine it. Guided Local Search (GLS) is an advanced metaheuristic that enhances local search by introducing a guidance mechanism to escape local optima [72, 66, 62]. GLS achieves this by modifying the objective function with penalties on certain solution features that appear in locally optimal solutions [73]. For the Traveling Salesperson Problem (TSP), the most natural features to penalize are the edges of the tour [74].
@@
HiFo-Prompt automates the discovery of the update heuristic $H_{\text{update}}$. Instead of relying on a static, human-designed rule, we task a LLM with synthesizing a complete, executable Python function to serve as $H_{\text{update}}$.
@@ B.2 Online Bin Packing Problem
The objective is to minimize the total number of bins used after all $T$ items have been placed [77, 78]. If we let $y_{j}=1$ if bin $j$ is used and $y_{j}=0$ otherwise, the goal is to minimize $\sum_{j}y_{j}$ [79].
@@
where $\sigma_{j}$ represents the "desirability" of placing item $a_{t}$ into bin $B_{j}$. The placement policy is then to assign the item to the valid bin with the highest score:
$j^{*}=\underset{j\,|\,c_{j}\geq s_{t}}{\arg\max}\,\sigma_{j}$ (19)
@@ B.3 Flow Shop Scheduling Problem
Due to the problem’s complexity, we employ a local search metaheuristic [84, 9]. To prevent the search from being trapped in local optima, we use the HiFo-Prompt framework to design a sophisticated guidance strategy automatically. When the search converges to a locally optimal sequence $\pi^{*}$, HiFo-Prompt-generated heuristic, $H_{\text{guide}}$, is invoked. It takes the current sequence, the original processing time matrix, and problem dimensions to produce both a new time matrix $T^{\prime}$ and a designated list of jobs to perturb, $J_{\text{perturb}}$:
@@ B.4 Bayesian Optimization
Bayesian Optimization (BO) [50] and its ongoing development are of paramount importance, as it provides the leading framework for sample-efficiently navigating the complex, high-cost search spaces prevalent in modern science and engineering. Bayesian Optimization (BO) has emerged as a principal framework for this task, excelling in applications like hyperparameter tuning and automated scientific discovery. The power of BO lies in its sample efficiency. It builds a probabilistic surrogate model of the objective function. It then uses an acquisition function [85, 86] to intelligently decide where to sample next, thereby minimizing the number of costly evaluations.
Our work addresses a particularly demanding variant: cost-aware BO [87, 88, 15]. In this setting, each function evaluation $f(\mathbf{x})$ has a heterogeneous and unknown cost, denoted by $c(\mathbf{x})$. The goal is to find the global maximum of $f(\mathbf{x})$ within a fixed total budget $B_{\text{total}}$. This requires sequentially choosing evaluation points $\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ to maximize the final outcome, subject to the constraint that $\sum_{i=1}^{N}c(\mathbf{x}_{i})\leq B_{\text{total}}$ [89]. To manage this, the BO agent maintains two surrogate models, typically Gaussian Processes (GPs) [90]. One GP models the objective function, predicting its posterior mean $\mu_{f}(\mathbf{x})$ and standard deviation $\sigma_{f}(\mathbf{x})$. A second GP models the evaluation cost, providing a cost prediction $\mu_{c}(\mathbf{x})$.
@@ C.1 Traveling Salesman Problem
To demonstrate the generality and robustness of the heuristic designed by our proposed method, we conduct a comprehensive evaluation using the real-world benchmark dataset TSPLib [59]. TSPLib is a well-established collection of the TSP instances, widely used in the research community for benchmarking optimization algorithms. For our experiments, we select a diverse subset of TSP instances from TSPLib, specifically focusing on those containing no more than 500 nodes. This selection criterion ensures a manageable problem scale while still retaining the complexity necessary to assess algorithmic performance effectively.
name EoH MCTS ReEvo Hsevo Ours
@@
mean 21.519% 17.629% 17.401% 19.168% 12.843%
Table 6: Results on TSPLib. The best results are highlighted in bold.
In our evaluation framework, we adopt a step-by-step construction approach to solve the TSP, which incrementally builds a tour by selecting the next node based on a learned heuristic. This paradigm allows us to evaluate the quality of decisions made at each step and better observe the contribution of the designed heuristic to the final solution quality. To rigorously assess the effectiveness and competitiveness of our proposed heuristic, we compare its performance against state-of-the-art LLM-based AHD baselines. These methods represent recent advances in leveraging large language models for combinatorial optimization tasks and serve as strong comparative baselines in our study. The experimental results, which include performance metrics, are summarized in Table 6. These results provide empirical evidence that our method not only performs competitively but also generalizes well across a variety of TSP instances in real-world scenarios.
@@ C.2 Online Bin Packing Problem
For each combination of bin capacity and problem size, the dataset includes five unique instances. Performance is quantified by the average percentage gap to the known theoretical lower bound across these instances, where a smaller value signifies a more efficient and effective packing solution. The empirical results unequivocally demonstrate the superior performance of the heuristic we have discovered. As shown in the table, our method consistently outperforms all baseline methods in nearly every setting, securing the smallest average gap in 8 out of the 9 distinct configurations tested. This highlights a remarkable level of consistency and dominance. The only exception is the (Capacity=100, Size=10k) case, where Funsearch achieves a marginally lower gap. However, our method’s strong performance across the entire parameter space underscores its greater reliability and generalizability compared to methods that may excel only in specific, narrow scenarios. Notably, our approach scales gracefully, achieving extremely low gap values (e.g., 0.41%, 0.42%) on the largest and most complex problems, significantly surpassing both traditional algorithms and other state-of-the-art LLM-driven frameworks. This comprehensive evaluation on challenging instances further validates the efficacy of our evolutionary framework, showcasing its capacity to discover sophisticated and high-performance heuristics that are robust across varied problem characteristics.
Capacity Size First Fit* Best Fit* Funsearch* EoH ReEvo MCTS-AHD HSEvo Ours
100 1k 5.32% 4.87% 3.78% 4.43% 3.63% 3.38% 3.53% 1.99%
5k 4.40% 4.08% 0.80% 1.07% 0.78% 0.99% 0.79% 0.66%
10k 4.44% 4.09% 0.33% 0.71% 0.35% 0.84% 0.34% 0.44%
300 1k 4.93% 4.48% 4.93% 4.48% 7.34% 3.21% 8.03% 2.01%
5k 4.18% 3.83% 1.07% 1.07% 4.47% 0.95% 5.41% 0.63%
10k 4.20% 3.87% 0.49% 0.73% 4.05% 0.85% 5.01% 0.41%
500 1k 4.97% 4.50% 6.75% 4.50% 5.92% 3.20% 6.45% 1.95%
5k 4.27% 3.91% 1.47% 1.06% 3.24% 0.95% 4.04% 0.66%
10k 4.28% 3.95% 0.74% 0.70% 2.79% 0.85% 3.49% 0.42%
@@ C.3 Flow Shop Scheduling Problem
Gap Obj. Gap Obj.
20 5 0.21% 1225.1 -0.01% 1222.4
10 0.24% 1517.4 0.15% 1515.9
20 0.18% 2239.1 0.10% 2237.4
50 5 0.02% 2736.8 0.00% 2736.4
10 0.29% 3006.2 0.12% 3001.2
20 0.75% 3759.1 0.60% 3753.4
100 5 -0.04% 5244.8 -0.05% 5244.5
10 0.17% 5639.8 0.09% 5635.3
20 0.93% 6421.6 0.47% 6392.6
200 10 0.25% 10699.3 0.08% 10681
20 0.94% 11480.7 0.63% 11444.9
Table 8: Results on FSSP. The best results are highlighted in bold.

### v2

@@ A.3 Knowledge-Augmented Evolutionary Computation
To extend the capabilities of evolutionary algorithms beyond simple adaptive control, a prominent research direction involves incorporating an explicit knowledge component. Cultural Algorithms [9, 62] provide a classic framework for this idea, decoupling the evolutionary system into two interacting spaces: a population space containing candidate solutions $P^{(t)}$ and a belief space $K_{t}$ serving as a repository of experiential knowledge [32]. These two spaces co-evolve through a dual-inheritance communication protocol.
@@
Currently, the successes of the population are fed back into the belief space. An acceptance function $\mathcal{A}$ identifies and extracts potentially valuable experiences, $\mathcal{A}(P^{(t)})$, from the current population. These are then used by a knowledge update function $\mathcal{U}_{K}$ to update the belief space:
@@ B.1 Traveling Salesman Problem
The Traveling Salesman Problem (TSP) [35, 58] is a canonical NP-hard combinatorial optimization problem. Given a set of $N$ cities and the distances between each pair, the objective is to find the shortest possible tour that visits each city exactly once before returning to the starting city.
@@
Regarding the integration of our method with the framework, we will discuss this in more detail. While fast, the myopic nature of such simple rules often leads to suboptimal solutions. Our approach, HiFo-Prompt, addresses this limitation by automating the discovery of more sophisticated construction heuristics. We retain the foundational step-by-step framework where a solution is built incrementally.
During training phase, we replace the fixed, handcrafted decision logic with a heuristic function synthesized by an LLM within our evolutionary framework. Crucially, rather than employing the LLM for online, per-step inference during the solving process, our method’s core contribution lies in the offline synthesis of a complete, executable heuristic function.
During inference phase, this generated function—not the LLM itself—that is invoked at each step $t$. This function receives the full problem state, including the current partial tour $S_{t}$, the set of unvisited candidate nodes $U_{t}$, and the global distance matrix $C$ [29, 28], and computes the next node to add to the tour locally and efficiently.
@@
Improvement heuristics, such as local search, start with a complete tour and iteratively refine it. Guided Local Search (GLS) is an advanced metaheuristic that enhances local search by introducing a guidance mechanism to escape local optima [58, 57]. GLS achieves this by modifying the objective function with penalties on certain solution features that appear in locally optimal solutions [61]. For the Traveling Salesperson Problem (TSP), the most natural features to penalize are the edges of the tour [52].
@@
HiFo-Prompt automates the discovery of the update heuristic $H_{\text{update}}$. Instead of relying on a static, human-designed rule, we task an LLM with synthesizing a complete, executable Python function to serve as $H_{\text{update}}$.
@@ B.2 Online Bin Packing Problem
The objective is to minimize the total number of bins used after placing all $T$ items [15, 65]. If we let $y_{j}=1$ if bin $j$ is used and $y_{j}=0$ otherwise, the goal is to minimize $\sum_{j}y_{j}$ [45].
@@
where $\sigma_{j}$ represents the desirability of placing item $a_{t}$ into bin $B_{j}$. The placement policy is then to assign the item to the valid bin with the highest score:
$j^{*}=\underset{\{j\,|\,c_{j}\geq s_{t}\}}{\arg\max}\,\sigma_{j}$ (19)
@@ B.3 Flow Shop Scheduling Problem
Due to the problem’s complexity, we employ a local search metaheuristic [29]. To prevent the search from being trapped in local optima, we use the HiFo-Prompt framework to design a sophisticated guidance strategy automatically. When the search converges to a locally optimal sequence $\pi^{*}$, HiFo-Prompt-generated heuristic, $H_{\text{guide}}$, is invoked. It takes the current sequence, the original processing time matrix, and problem dimensions to produce both a new time matrix $T^{\prime}$ and a designated list of jobs to perturb, $J_{\text{perturb}}$:
@@ B.4 Bayesian Optimization
Bayesian Optimization (BO) [46] and its ongoing development are of paramount importance, as it provides the leading framework for sample-efficiently navigating the complex, high-cost search spaces prevalent in modern science and engineering. Bayesian Optimization (BO) has emerged as a principal framework for this task, excelling in applications like hyperparameter tuning and automated scientific discovery. The power of BO lies in its sample efficiency. It builds a probabilistic surrogate model of the objective function. It then uses an acquisition function [24] to intelligently decide where to sample next, thereby minimizing the number of costly evaluations.
Our work addresses a particularly demanding variant: cost-aware BO [49, 64, 67]. In this setting, each function evaluation $f(\mathbf{x})$ has a heterogeneous and unknown cost, denoted by $c(\mathbf{x})$. The goal is to find the global maximum of $f(\mathbf{x})$ within a fixed total budget $B_{\text{total}}$. This requires sequentially choosing evaluation points $\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ to maximize the outcome, subject to the constraint that $\sum\limits_{i=1}^{N}c(\mathbf{x}_{i})\leq B_{\text{total}}$ [25]. To manage this, the BO agent maintains two surrogate models, typically Gaussian Processes (GPs) [39]. One GP models the objective function, predicting its posterior mean $\mu_{f}(\mathbf{x})$ and standard deviation $\sigma_{f}(\mathbf{x})$. A second GP models the evaluation cost, providing a cost prediction $\mu_{c}(\mathbf{x})$.
@@ C.1 Traveling Salesman Problem
To demonstrate the generality and robustness of the heuristic designed by our proposed method, we conduct a comprehensive evaluation using the real-world benchmark dataset TSPLib [40]. TSPLib is a well-established collection of TSP instances, widely used in the research community for benchmarking optimization algorithms.
For our experiments, we select a diverse subset of TSP instances from TSPLib, specifically focusing on those containing no more than 500 nodes. This selection criterion ensures a manageable problem scale while still retaining the complexity necessary to assess algorithmic performance effectively.
In our evaluation framework, we adopt a step-by-step construction approach to solve the TSP, which incrementally builds a tour by selecting the next node based on a learned heuristic. This paradigm allows us to evaluate the quality of decisions made at each step and better observe the contribution of the designed heuristic to the final solution quality.
To rigorously assess the effectiveness and competitiveness of our proposed heuristic, we compare its performance against state-of-the-art LLM-based AHD baselines. These methods represent recent advances in leveraging large language models for combinatorial optimization tasks and serve as strong comparative baselines in our study. The experimental results, which include performance metrics, are summarized in Table 6. These results provide empirical evidence that our method not only performs competitively but also generalizes well across a variety of TSP instances in real-world scenarios. In addition, we further conduct experiments on small-scale instances (TSP10 and TSP20) under the step-by-step construction framework for TSP. The results are presented in Table 7.
Table 6: Results on TSPLib. The best results are highlighted in bold.
name EoH MCTS-AHD ReEvo HSEvo Ours
@@
avg. 21.519% 17.629% 17.401% 19.168% 12.843%
Table 7: Results on TSP with step-by-step construction in 10 and 20.
Method TSP10 TSP20
Gap Time(s) Gap Time(s)
LKH3 0.000% 6.492 0.000% 24.948
POMO 0.246% - 0.248% -
LEHD 0.183% - 0.010% -
EoH 7.148% 0.042 10.064% 0.168
ReEvo 5.227% 0.228 6.811% 1.215
HSEvo 5.461% 0.689 7.950% 3.273
MCTS-AHD 4.829% 0.440 8.045% 4.087
Ours 1.654% 0.709 3.619% 12.961
@@ C.2 Online Bin Packing Problem
For each combination of bin capacity and problem size, the dataset includes five unique instances. Performance is quantified by the average percentage gap to the known theoretical lower bound across these instances, where a smaller value signifies a more efficient and effective packing solution.
The empirical results demonstrate the superior performance of the heuristic we have discovered. As shown in the table, our method consistently outperforms all baseline methods in nearly every setting, securing the smallest average gap in 8 out of the 9 distinct configurations tested. This highlights a remarkable level of consistency and dominance. The only exception is the (Capacity=100, Size=10k) case, where Funsearch achieves a marginally lower gap.
However, our method’s strong performance across the entire parameter space underscores its greater reliability and generalizability compared to methods that may excel only in specific, narrow scenarios. Notably, our approach scales gracefully, achieving extremely low gap values (e.g., 0.41%, 0.42%) on the largest and most complex problems, significantly surpassing both traditional algorithms and other state-of-the-art LLM-driven frameworks. This comprehensive evaluation on challenging instances further validates the efficacy of our evolutionary framework, showcasing its capacity to discover sophisticated and high-performance heuristics that are robust across varied problem characteristics.
@@
Capacity Size First Fit Best Fit EoH ReEvo HSEvo MCTS-AHD Ours
100 1k 5.32% 4.87% 3.10% 3.63% 3.51% 3.38% 2.19%
5k 4.40% 4.08% 1.02% 0.78% 1.91% 0.99% 0.69%
10k 4.44% 4.09% 0.80% 0.35% 1.68% 0.84% 0.42%
300 1k 4.93% 4.48% 3.04% 7.34% 6.88% 3.21% 2.08%
5k 4.18% 3.83% 1.00% 4.47% 5.47% 0.95% 0.66%
10k 4.20% 3.87% 0.78% 4.05% 5.26% 0.85% 0.39%
500 1k 4.97% 4.50% 3.04% 5.92% 5.80% 3.20% 2.07%
5k 4.27% 3.91% 1.00% 3.24% 4.39% 0.95% 0.66%
10k 4.28% 3.95% 0.78% 2.79% 4.14% 0.85% 0.40%
@@ C.3 Flow Shop Scheduling Problem
Table 9: Results on FSSP. The best results are highlighted in bold.
@@
Gap Time(s) Gap Time(s)
20 5 0.25% 7.5 -0.01% 5.5
10 0.31% 13.0 0.17% 8.8
20 0.20% 23.8 0.10% 18.5
50 5 0.01% 45.1 0.00% 29.7
10 0.29% 83.9 0.17% 50.6
20 0.84% 168.4 0.58% 101.1
100 5 -0.02% 230.2 -0.04% 146.7
10 0.23% 299.6 0.13% 243.5
20 0.94% 305.2 0.51% 303.0
200 10 0.37% 337.1 0.12% 317.2
20 1.23% 334.6 0.71% 321.4